OpenStudy (anonymous):

Write in standard form of parabola: 1. 9x^2+36x+25y^2-50y=164 2. x^2+y^2-10x+4y+27=0 3. 45y^2-320x^2+6=2886

OpenStudy (imstuck):

Do you have any ideas on this?

OpenStudy (anonymous):

@IMStuck This would basically be simplifying it right?

OpenStudy (anonymous):

no since \( x^2+y^2-10x+4y+27=0\) is not a parabola!

OpenStudy (anonymous):

actually, none are parabolas!

OpenStudy (anonymous):

Well I just have to write all of them in standard form starting from finding the vertex

OpenStudy (anonymous):

Wow @satellite73 , you are clearly very knowledgeable :)

OpenStudy (anonymous):

you cannot find the vertex they are not parabolas

OpenStudy (anonymous):

Ok well how do i write these in standard form?

OpenStudy (anonymous):

\[9x^2+36x+25y^2-50y=164\]is an ellipse

OpenStudy (imstuck):

I was going to say that these are not parabolas at all. The first one could be a hyperbola. It most likely is!

OpenStudy (anonymous):

you can write it in standard form of an ellipse if you like you will end up with\[\frac{(xh)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (imstuck):

No forget that...ellipse!

OpenStudy (anonymous):

Ok I am still very confused... can you demonstrate one of them?

OpenStudy (anonymous):

\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]rather it is an annoying piece of algebra called "completing the square"

OpenStudy (imstuck):

i personally like completing the square! No?

OpenStudy (anonymous):

i can walk you through the first one if you like \[9x^2+36x+25y^2-50y=164\]start by factoring out the common factor of the \(x\) and \(y\) terms and start with \[9(x^2+4x)+25(y^2-2y)=164\]

OpenStudy (anonymous):

then take half of the coefficient of the linear term, write as a complete square via \[25(x+2)^2+25(y-1)^2=164+9\times 2^2+25\times 1^2\]

OpenStudy (imstuck):

You are really doing a good job for madisonhope96. Yay for you. These are not the easiest things to get through!

OpenStudy (anonymous):

ok i follow you

OpenStudy (anonymous):

damn typo!!\[9(x+2)^2+25(y-1)^2=164+9\times 2^2+25\times 1^2\]

OpenStudy (anonymous):

Ok, I still got what you meant from the first demonstration! :)

OpenStudy (anonymous):

arithmetic on the right gives \[9(x+2)^2+25(y-1)^2=225\]

OpenStudy (anonymous):

divide by \(225\) and get \[\frac{(x+1)^2}{25}+\frac{(y-1)^2}{9}=1\]

OpenStudy (anonymous):

which looks like \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]exactly

OpenStudy (imstuck):

I'm getting that the second one is a circle.

OpenStudy (anonymous):

Thank you so much! Now how would you find the minor vertices of that problem if you don't mind me asking?

OpenStudy (anonymous):

yet another typo!\[\frac{(x+2)^2}{25}+\frac{(y-1)^2}{9}=1\]

OpenStudy (anonymous):

center is \((-2,1)\)

OpenStudy (anonymous):

So that is the only minor vertices? Im so confused sorry

OpenStudy (anonymous):

minor axis vertices*

OpenStudy (anonymous):

i have never heard of "minor axis vertices"

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

if you want the length of the major axis, it is \(10\) and the length of the minor axis is \(6\)

OpenStudy (anonymous):

\[\frac{(x+2)^2}{25}+\frac{(y-1)^2}{9}=1\]\[\frac{(x+2)^2}{5^2}+\frac{(y-1)^2}{3^2}=1\]\[a=5,b=3\]

OpenStudy (imstuck):

The minor axis vertices lie on the axis that the ellipse is NOT centered about. In other words, if the x axis is the major vertice axis and the foci lie on this axis, then the other axis is the minor. In an ellipse, a is ALWAYS greater than b, and a = 5 and b = 3. The 5 is under the x, so the x is the major vertice and y is the minor.

OpenStudy (anonymous):

actually although it will not tell you how to get it, any information about this ellipse can be found here http://www.wolframalpha.com/input/?i=ellipse+%28x%2B2%29^2%2F25%2B%28y-1%29^2%2F9%3D1&dataset=&equal=Submit

OpenStudy (imstuck):

The coordinates for the major vertices are (3, 1) and (-7, 1) and the coordinates for the minor vertices are (-2, 4) and (-2, -2)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (imstuck):

Do you understand how to find the major and minor vertices? It's quite simple, just takes some explaining.

OpenStudy (anonymous):

Yes, that clarified a lot! Now the last thing I'm confused on is how to find the value of p for -2x^2+16x+24y-224=0???

OpenStudy (imstuck):

Ok so if you have a p to find, this tells me right away that this is a parabola.

OpenStudy (anonymous):

Ok, so I'm just confused as to where exactly I would start?

OpenStudy (imstuck):

You must complete the square in this one too. The standard form for a parabola is x^2=4py, where p is the distance that the focal point is from the vertex.

OpenStudy (imstuck):

i will do step by step as did mr. satellite73. Ok?

OpenStudy (anonymous):

Ok thank you!

OpenStudy (imstuck):

\[-2x ^{2}+16x=-24y+224\]Or\[2x ^{2}-16x=24y-224\]Complete the square on the x terms (this is an x^2 parabola, meaning it opens either upwards or downwards. I tell my students it is either a cup or a mountain).

OpenStudy (imstuck):

\[2(x ^{2}-8x)=24y-224\]\[2(x ^{2}-8x+16)=24y-224+2(16)\]\[2(x-4)^{2}=24y-192\]\[(x-4)^{2}=12y-96\]\[(x-4)^{2}=12(y-8)\]

OpenStudy (anonymous):

Ok I understand that better now

OpenStudy (imstuck):

Now here is the point at which you are able to solve for p. The method used for this is 4p=coefficient of y. Which is 4p = 12, so p = 3. That means that the focus is 3 units up from the vertex of (4, 8) and it is a "cup" parabola, because the focus is above the vertex. That's the whole equation; you kind of needed to do the whole thing in order to find p.

OpenStudy (imstuck):

These are not easy; it was THE hardest chapter for my Algebra 2 kids this past year. Keep at it, memorize the rules for each type of quadratic equation and it will make it easier for you. Promise.

OpenStudy (anonymous):

Oh! I originally got that, but thought I needed to do more at first! So p=3 is all i need? Thank you sooooo much!

OpenStudy (imstuck):

You've most very welcome!!!